A 25.00 mL solution of fruit juice containing vitamin C (C6H8O6) is reacted with 7.83 ml of a 0.5871 M solution of iodine (I2) according to the following equation:
a. balance the chemical equation
b. determine how many miligrams of vitamins C (C6H8O6) are in the fruit juice
c. what is the molarity of vitamin c in the fruit juice?
1
Expert's answer
2019-12-03T08:24:18-0500

(a)"C_6H_8O_6 + I_2 = C_6H_6O_6 + 2 H^{+} + 2 I^{-}"

b)no of moles of "I_2=0.5871\u00d77.83=4.6milli moles"

no of moles of "C_6H_8O_6=4.6milli moles"

Mass of "C_6H_8O_6=4.6\u00d7176.12=809.62 mg"

c)molarity=4.6/32.83=0.14M